40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 210 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  6522   Wed Apr 11 18:19:51 2012 JenneUpdateIOOMode matching recollections and conclusions

Another histogram.  This one allows the MMT mirror positions to move, the MMT incident angle for both curved optics to change, and the MC waist size and position to change.  The error quoted for the MC waist size measurement from 2010

was +\- 0.01mm, and the MC waist position was +\- 28mm.  

This histogram is showing that we're pretty sensitive to the MC waist measurement, which is used to define the beam.  We can be up to ~2% off in our ideal mode matching to the arms if we're using the incorrect initial beam for the telescope design.

ModeMismatch_MMTpos_MMTang_MCw0_MCwz_change_LowRes.png

  6521   Wed Apr 11 17:56:26 2012 JenneUpdateGeneralSummary of things to figure out with the IFO

Questions:

Power recycling gain

   * It should be ~40, but we observe/measure it to be ~7.  Even if mode matching of ~50% is assumed, gain is calculated to be ~15

   * Would like to measure PR gain independent of mode matching, if possible

Power recycling cavity mode matching

   * Reflectivity of PRMI was measured to be ~50%.  That's pretty high.  What's going on?

   * Even if we're mode matched to the arm, are we appropriately mode matched to the PRC?

Is beam from MC clipped in the Faraday?

   * We had to use MC axis for input pointing since PZTs aren't totally working.

   * Need to measure IPPOS beam for different MC alignments to see if horizontal waist measurement stays constant.

PRM flipped?

   * Not likely, but it can't hurt to confirm for sure.

   * Want to know, since it could give us a different plan for MMT moving than if the PRM is correct.

Thick optic non-normal incidence in IPPO - does this exaggerate astigmatism, which would help explain IPPOS measurement?

Is PRC waist same size / position as arm cavity waist, given the current "known" positions of all the optics?

   * How is this effected by moving the PRM?

------------------------------------------------------------------------------------

Measurements to take and what information they give us:

IPPOS beam scan, with MC as-is

   * Confirm (or not) IPPOS measurements from last week

IPPOS beam scan with different MC alignments

   * Will tell us about Faraday clipping, if any

AS beam scan, misaligned PRM, misaligned SRM, misaligned ITMX, single bounce from ITMY

   * Can only take this measurement if beam is bright enough, so we'll just have to try

   * Will confirm IPPOS measurement, but includes going through the thick PRM, so can compare to calculated intra-PRC mode

REFL beam scan (already done....is the data satisfactory? If so, no need to redo), single bounce off of PRM

   * Will tell us about the potential PRM flipping

   * Need to compare with calculated mode at REFL port for flipped or non-flipped PRM

Look at POP camera, see 2nd pass through cavity

   * Try to match 1st and 2nd pass.  If they don't match, we're not well matched to PRC mode

Look at beam directly on ETMY cage, then beam from ETM, bounce off ITM, back to ETM cage

   * If the beams are the same size, we're well matched to arm cavity mode

   * Use fancy new frame-grabber.

----------------------------------------------------------------------------

MMT code things to calculate, and what information it gives us:

REFL beam path, for PRM flipping comparison

Thick IPPO non-normal incidence - I'm not sure how to do this yet, since I only know how non-normal incidence changes effective radii of curvature, and this is a flat optic, so *cos(theta) or /cos(theta)  won't do anything to an infinite RoC

Compare PRC waist to arm cavity waist, using "known" optic positions

Mode matching sensitivity to MC waist measurements

Mode matching sensitivity to PRM position

  6520   Wed Apr 11 16:33:16 2012 JenneBureaucracyGeneral40m Meeting Action Items

Action Items from Last Week:

(non-existant)

-----------------------------------------------------------------------

Action Items this Week and LEAD PERSON:

Assemble and ship 4 TTs from LHO - SURESH

Prepare electronics for TTs (coil drivers) - JAMIE

In-air TT testing to confirm we can control / move TTs before we vent (starting in ~2 weeks) - SURESH

Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH

OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN

Static-only OAF noise budget (Adaptive noise budget as next step) - DEN

Black glass: big baffle pieces to clean&bake, get small pieces from Bob, put into baskets, make new basket for 1" pieces, get to clean&bake - KOJI

IPPOS beam measurement - SURESH with JENNE

AS beam measurement (if beam is bright enough) - SURESH and JENNE

Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE

Summary of IFO questions, measurements to take, and game plan - JENNE

Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH

  6519   Wed Apr 11 14:34:48 2012 JenneUpdateEnvironmentPlumbing guys are here

Quote:

I don't know why, but they're looking around on the roof, and inside our ceiling above the bathrooms.

 Bob tells me that the carpenter is going to move the nitrogen bottles to the other side of the outside door, so that the plumbers can install a safety shower / eyewash right outside our door.

Also, the carpenter just mounted a new glass door cabinet from Bob's lab in the IFO room, so we have some new storage space.

  6518   Wed Apr 11 12:25:11 2012 RyanUpdateComputersUpdating aLIGO Conlog

Over the next few days, I will be working on upgrading the aLIGO Conlog install to include new bugfixes distributed by Patrick T.  The currently running conlog *should* not be affected, but please let me know if it is (ryan.fisher@ligo.org).

  6517   Tue Apr 10 23:56:44 2012 ranaUpdateComputersSensoray

Now that Mike has got the Sensoray working, Jenne/Suresh should grab some new images of the ETM cage as Keiko did so that we can analyze them for another mode matching diagnostic.

  6516   Tue Apr 10 17:02:29 2012 JenneUpdateIOOMode matching recollections and conclusions

...Mostly just recollections at this point.

I re-looked at the mode matching's sensitivity to misplaced optics.  Here is the plot that the original MMT code from 2010 spits out:

IMMT_positionSensitivity_plotted10Apr2012_LowRes.png

What this plot is telling us is that we should lose no more than 0.1% of mode matching "goodness" if we messed up the curved optic's positions by up to 2 cm.  If we can't place optics to within 2 cm, we might as well go back to optics kindergarten, because that's pretty lame.

UPDATE: Here is a histogram using the new code, which definitely includes the non-unity index of refraction for the transmissive optics and the Faraday.  The only optics which are permitted to move are the 2 curved optics, and they are allowed a stdev of 20mm.  Again, we shouldn't be doing worse than ~99% mode matching, even if we're 2cm off from the MMT positions that we measured with a ruler.  This histogram only has 300 iterations, since it takes quite a while (~0.5sec) to calculate each iteration.  Note this is mode overlap using the measured MC waist, propagated through optics, compared to the ideal arm mode.  This is completely ignoring the IPPOS measurements so far.

Histogram_20mmStDev_MMTmovesOnly_LowRes.png

 

UPDATE 2: Allowed 5 degrees of incident angle motion for both curved optics, which changes the astigmatism of the beam downstream.  Still, no big change from ~99% mode matching efficiency.  Again, this doesn't include any information from the IPPOS measurements.  3000 iterations this time around, since I didn't need my computer.  Curved optics still allowed to move back and forth by 2cm. 

Histogram_20mmStDev_MMTmovesAndTilts_LowRes.png 

More meditations and conclusions to follow...  currently running hist code to allow tilt of optics, to account for astigmatism changes also. 

Suresh and I are going to do some beam measurements tomorrow with the beamscanner, and then we will do a few measurements with the razor blade technique, to confirm that we're doing things okey dokey.

  6515   Tue Apr 10 13:42:54 2012 JenneUpdateEnvironmentPlumbing guys are here

I don't know why, but they're looking around on the roof, and inside our ceiling above the bathrooms.

  6514   Tue Apr 10 11:08:29 2012 taraUpdatePSLcurved mirror behind AOM removed

We removed the curved mirror behind the AOM (ROC=0.3m) on PSL table. The mirror is now in PSL lab. See PSL:905 for more detail.

  6513   Mon Apr 9 20:02:19 2012 Mike J.UpdateComputersSensoray

The highest resolution available is 720x480 pixels. Bit depth of captured images and video is most likely 16 bits per pixel. Video may be captured raw as well, which will be necessary for image subtraction/enhancement, however it cannot currently be played raw. A captured image is shown below, along with MP4 video.

out_0.jpg

 

  6512   Mon Apr 9 18:18:14 2012 kiwamuUpdateLSCDRMI time series

Here is a time series when the DRMI is being locked.

You can see that the AS110 goes up because the SRCL is engaged and amplifies the 55 MHz sidebands.

 time_series1.png

  6511   Mon Apr 9 17:03:38 2012 DenUpdateEnvironmentLms vs Wiener

I tried to figure out why offline LMS filter subtract seismic noise much better from MC_F then the Wiener filter. I did the calculations twice - with my codes and with Matlab in-build functions, the results are the same. So this is not a code error.

The coherence between GUR 1, 2 and MC_F is still poor. Wiener filter is linear and its performance is confined to the frequency ranges where we see coherence. Lms filter is non-linear and it may be possible to subtract the noise even if non-linear effects are present in the system.

gur12_mcl.png

I've checked seismometer readout box again. I've soldered 50 Ohms to plus and minus inputs to VERT 1,2 N/S 1,2, E/W 1,2 - GUR 1 and 2 use these channels. Then I put the box back and connected it to the ADC.

seismboxnoise.png

The plot shows that the readout box noise is below the ADC noise. It is possible that amplifiers introduce non-linear effects. To check this I plotted the coherence between OSEM sensors and GUR1X signal:

gur1_osem.png

The coherence between OSEM sensors and GUR1X is pretty good, so may be witness path is not responsible for low coherence at 0.1 - 0.5 Hz between MC_F and GUR 1,2. IT seems that MC_F is bad at low frequencies. I terminated the input to the Channel 1 of the Pentek Generic board, where MC_F is plugged in.

mcl_noise.jpg

ADC is also good. Something else is wrong.

  6510   Mon Apr 9 15:09:34 2012 JenneUpdateLSCLocked MICH

Quote:

I was going to try some locking, but things are a little too noisy. 

Just so Kiwamu knows what I did today, in case he comes back....

I ran LSCoffsets, and aligned both X and Y arms and saved their positions, and aligned MICH, and saved the BS position. 

I'll play with it more later, when there aren't trucks driving around outside that I can hear / feel in the control room.

 After giving up on locking, the MC is getting unlocked every now and again (2 times so far in the last few minutes) from transient seismic stuff.

  6509   Mon Apr 9 15:02:30 2012 JenneUpdateLSCLocked MICH

I was going to try some locking, but things are a little too noisy. 

Just so Kiwamu knows what I did today, in case he comes back....

I ran LSCoffsets, and aligned both X and Y arms and saved their positions, and aligned MICH, and saved the BS position. 

I'll play with it more later, when there aren't trucks driving around outside that I can hear / feel in the control room.

  6508   Sat Apr 7 06:58:34 2012 kiwamuUpdateLSCDRMI lock : lost good alignment

Somehow I lost the good alignment, where the lock can be frequently acquired and hence I didn't go further ahead.

I will try locking the DRMI during the weekend again. My goal is to take time series when the DRMI is being locked and sensing matrix.

Quote from #6506

Currently SRM is slightly misaligned such that the MICH optical gain at AS55Q doesn't increase so much with the presence of SRM.

With this condition I was able to acquire the lock more frequently than how it used to be on the Wednesday.

The next step is to gradually align SRM, to optimize the controls and to repeat this process several times until SRM is fully aligned.

 

  6507   Sat Apr 7 02:01:29 2012 Mike J.UpdateComputersProjector Cable Management

I replaced the projector video and power cables with longer ones, and zip-tied them to the ceiling and wall so they don't block the image.

projector_cables.jpg

  6506   Sat Apr 7 01:56:05 2012 kiwamuUpdateLSCOSA signal in DRMI condition

It wasn't a dream or illusion -- I was locking the DRMI to the right condition last Wednesday (#6489).

Here is a snap shot of the AS-OSA signal taken today when the DRMI was locked with the same control settings (#6489).

The blue curve is data taken when the PRMI was locked for comparison.

You can see that both the upper and lower 55 MHz sideband are amplified by the SRC.

OSA.png

 

(Some notes)

Currently SRM is slightly misaligned such that the MICH optical gain at AS55Q doesn't increase so much with the presence of SRM.

With this condition I was able to acquire the lock more frequently than how it used to be on the Wednesday.

The next step is to gradually align SRM, to optimize the controls and to repeat this process several times until SRM is fully aligned.

Quote from #6489

A first thing I tried was : flipping the control sign of the SRCL while keeping the same control setups for the PRCL and MICH.

Occasionally the DRMI was "sort of" locked and hence I believe this setup must be a good starting point.

  6505   Sat Apr 7 01:45:02 2012 Mike J.UpdateComputersEven Better Hysteresis Model and Plots

 The new hysteresis model is slightly based on the SHO equation, but with the force being out of phase with the position by an amount of hysteresis {x(t)=Amp*sin(freq*t), F(t)=Amp*sin(freq*t+Hyst)}. The new model can be found at /users/mjenson/matlab/hyst_v_3.mdl.  Pictures are: new hysteresis model, x(t) subsystem in new model[xh''(t) only lacks -1 multiplier and includes hysteresis variable], new plots.

 hyst_v_3.pnghyst_v_3-x(t).pnghyst_v3.png

  6504   Sat Apr 7 00:31:12 2012 keikoUpdateLSCRAM simulation for Full ifo

I didn't understand how CARM can be decreased 2 orderes of magnitude and PRCL can be INCREASED by such small offsets (see the matrix quoted).

Apparently it was because of an optical-spring ish effect from the "detuning" (which is actually RAM position offsets). I put two plots which are CARM and PRCL tranfer functions to REFL f1 or POP f1, when there is a slight PRCL offset (0, 1e-14m, and 1e-15m cases are plotted). Looking at these plots, it was not a good idea to calculate the LSC matrix in DC because they are affected by this detuning a lot. I'll try f = 150 Hz for the matrix.

plot4a.pngplot4b.png

Quote:

*** Iteration *** 

%*.*.*. Resulting matrix w/ RAM .*.*.*

REFL f1 : 0.039125    -0.000000    0.003665       0.000005    -0.000007 

  AS f2 : 0.000010    1.000431    0.000009       -0.003500    -0.000002 

 POP f1 : 156.420221    -0.000246    15.586838    0.019406    -0.000154 

 POP f2 : 1.255806    -0.154275    0.047313       1.000008    0.024285 

 POP f2 : -34.814720    -0.006600    -1.884850    0.042950    1.000000 

Offsets converged to:

PRCL =  2.1e-15, MICH = 1.1e-17, SRCL = -3.8e-15, CARM = 2.2e-16, DARM = 0  

  6503   Fri Apr 6 20:38:41 2012 Mike J.UpdateComputersSensoray

 Turns out that the "MPEG-4 VES" video format is just bad for captured video.  Everything except "MP4" and "MPEG-TS" works for streaming, and "MP4" and "MPEG-TS" seem to be the only captured formats that can be viewed properly.

  6502   Fri Apr 6 20:24:31 2012 Mike J.UpdateComputersSensoray

The Sensoray device is currently viewing Monitor 4 and plugged into Pianosa.  The user interface is run at /home/controls/Downloads/sdk_2253_1.2.2_linux/python demo.py. It can preview and capture the video stream, however the captured files are terrible. I believe it has something to do with the bitrate, since the captured video with lower bitrates are not as bad as the ones with higher bitrates, but  I am not certain.

  6501   Fri Apr 6 20:05:12 2012 JamieSummaryGeneralLaser Emergency Shutoff

We reset the interlock and restarted the PSL.  The end AUX lasers seem to have come back online fine.  PMC and mode cleaner locked back up quickly.

  6500   Fri Apr 6 19:40:57 2012 Mike J.SummaryGeneralLaser Emergency Shutoff

I accidently shut off the laser at 19:34 with the emergency shutoff button while trying to tap into a video line for the Sensoray device.

  6499   Fri Apr 6 19:04:35 2012 JenneUpdatePEMSTS releveled, GUR2 plugged in

[Den, Jenne]

We were wondering why the STS-2 signal was funny.  When I went to look at it, the X-axis indicator was pointing ~45deg from the x-axis, so that it was pointing between the arms of the IFO.  Also, the bubble in the level was totally stuck on one side.  We locked the masses, and I put the seismometer back to the correct orientation, and then leveled it.  We unlocked the masses and turned the power back on, and hit the auto-zero button a few times.  Right now the X-axis signal is fine, but Y and Z are still railed, but it's been like 24 seconds, not 24 hours since we last hit auto zero, so there's still some time to wait.

Also, GUR2 was unplugged on both ends of the cable.  We plugged it back in.  However, it looks like the *seismometer* labeled #1 is now plugged into *channels* GUR2, and the seismometer labeled #2 is plugged into channels GUR1.  Recall that Den has only modified X, Y, Z for GUR1 channels, not any other channels in the breakout box.

  6498   Fri Apr 6 16:35:37 2012 DenUpdateComputersc1ioo

c1ioo computer can not connect to the framebuilder and everything is red in the status for this machine, C1:FEC-33_CPU_METER is not moving.

EDIT by KI:

 We rebooted the c1ioo machine, but none of the ftont end model came back. It looked like they failed the burt process for some reasons according to dmesg.

Then we restarted each front end model one by one, and every time after immediately we restarted it we hit the 'BURT' button in the GDS screen.

Everyone came back to the normal operation.

  6497   Fri Apr 6 16:22:15 2012 DenUpdateEnvironmentseism box

I've changed R2 resistor in the seism box for the VERT 1 channel from 464 Ohm to 1051 Ohm to reduce the gain of this channel by a factor of 2. This should help the GUR1Z signal not to be corrupted inside the AA box, so we can use it in the adaptive filtering.

  6496   Fri Apr 6 15:06:05 2012 DenUpdateIOO1 Hz resonance

I think we can try to damp 1 Hz resonance more. In September it was not seen because of the digital noise. After we've figured it out, 1 Hz resonance began to be more clear (blue line).

psd_mcl.jpg

Now applying oaf we reduce the effect of the stack and the 1 Hz resonance is even more clear:

mcl.jpg

 

  6495   Fri Apr 6 14:39:21 2012 JamieUpdateComputersRAID array is rebuilding....

The RAID (JetStor SATA 416S) is indeed resyncing itself after a disk failure.  There is a hot spare, so it's stable for the moment.  But we need a replacement disk:

    RAID disks:  1000.2GB Hitachi HDT721010SLA360

Do we have spares?  If not we should probably buy some, if we can.  We want to try to keep a stock of the same model number.

Other notes:

The RAID has a web interface, but it was for some reason not connected.  I connected it to the martian network at 192.168.113.119.

Viewing the RAID event log on the web interface silences the alarm.

I retrieved the manual from Alex, and placed it in the COMPUTER MANUALS drawer in the filing cabinet.

  6494   Fri Apr 6 11:32:09 2012 JenneUpdateComputersRAID array is rebuilding....

Suresh reported to Den, who reported to me (although no elogs were made.....) that something was funny with the FB.  I went to look at it, and it's actually the RAID array rebuilding itself.  I have called in our guru, Jamie, to have a look-see.

  6493   Fri Apr 6 11:14:34 2012 JenneUpdateAdaptive Filteringstatic and adaptive

Quote:

I've run static and adaptive filters simultaneously. AA32 filters rotate the phase of the witness signals GUR1X and GUR1Y and now the performance of the static filter is worse. Next time I'll recalculate Wiener filter coefficients taking this into account. But still 2 filters together can deal with a stack better.

static_oaf.pdf

 This is super awesome!  I'm totally excited!!

  6492   Fri Apr 6 10:31:07 2012 DenUpdateAdaptive Filteringstatic and adaptive

I've run static and adaptive filters simultaneously. AA32 filters rotate the phase of the witness signals GUR1X and GUR1Y and now the performance of the static filter is worse. Next time I'll recalculate Wiener filter coefficients taking this into account. But still 2 filters together can deal with a stack better.

static_oaf.pdf

  6491   Fri Apr 6 09:57:24 2012 DenUpdateAdaptive Filteringstatic starts to work

I made static filter to work to evaluate the actuator TF.. Here is the result of static filtering:

static1-crop.pdf

 What I did:

 I did offline simulation of the MC_F Wiener filtering using 2 witness signals - GUR1X and GUR1Y. I've downsampled the data from 2048 to 128 Hz and applied the Wiener filter with 10000 for each witness channel:

wiener_filtering.pngcoeffs.png

                                            Result of the filtering                                                                                     Filter coefficients for gur1x and then gur1y

xTF.pngyTF.png

                                         Gur1x -> MC_F transfer function                                                                          Gur1y -> MC_F transfer function

Then using vectfit I approximated obtained transfer functions in the region 0.5 - 20 Hz. I used a window function and then weights to get a more precise result in this range using only 8 poles and zeros.

xfitting.pngyfitting.png

I obtained the zpk-model for each witness channel and entered it into the FOTON splitting it into 2 parts before that because FOTON does not like too long filters. These zpk-models are at the C1:OAF-STATIC_STATMTX_8_8 and C1:OAF-STATIC_STATMTX_8_9 filter banks.

GUR1X:

z =

  7.527339430029315 +31.603999069997801i
  7.527339430029230 -31.603999069997823i
 27.897703898191267 + 0.000000000000071i
 -6.437806394766186 + 9.893955654289517i
 -6.437806394766159 - 9.893955654289510i
  1.114401249545640 + 5.479278396987240i
  0.176877296954015 + 0.000000000000006i
  1.114401249545616 - 5.479278396987245i


p =

 -0.407251778925379 + 6.263247012022007i
 -0.407251778925379 - 6.263247012022007i
 -0.230672968859081 + 6.846868757063707i
 -0.230672968859081 - 6.846868757063707i
 -2.871419857491615 +13.707864827517826i
 -2.871419857491615 -13.707864827517826i
 -2.134260618362721 +18.319129999777648i
 -2.134260618362721 -18.319129999777648i


k =

     4.113285626223658e-04

GUR1Y

z =

 17.961416874092624 +13.631821252434328i
 17.961416874092642 -13.631821252434353i
 -8.788634771726304 + 7.653357335975781i
 -8.788634771726285 - 7.653357335975777i
 -0.037906973323273 + 5.133348020858679i
 -0.164348392996182 + 3.588803405511463i
 -0.164348392996187 - 3.588803405511474i
 -0.037906973323277 - 5.133348020858679i


p =

 -0.027577318242359 + 5.174655410828068i
 -0.027577318242359 - 5.174655410828068i
 -0.500384298611703 + 6.310552036591990i
 -0.500384298611703 - 6.310552036591990i
 -0.237055716999485 + 6.881204941979009i
 -0.237055716999485 - 6.881204941979009i
 -1.408223271160550 +14.874570175309771i
 -1.408223271160550 -14.874570175309771i


k =

    -2.723835471763049e-04

 Then I approximated the reversed actuator TF  and placed it to the C1:OAF-SUS_MC2_OUT filter bank. The gain to the static filter output is -1.

P.S. Also the static matrix was filled with 1 for some reason. Here is the script to fix it if if will be bad again

for i in {1..8}
do
    for j in {1..28}
    do
        element="C1:OAF-STATIC_STATMTX_"$i"_"$j"_GAIN"
        ezcawrite $element 0
    done
done

 

 

  6490   Thu Apr 5 18:24:55 2012 DenConfigurationAdaptive Filteringoaf starts to work

Today I tried to make the lms filter to work online. I played around with the signals (GUR1_X and MC_F) to pre-whiten them and in the end the following configuration worked out:

1. mu = 0.03, tau = 1e-5, downsample=8, nCoeff = 4000, delay = 5 (sample-and-hold delay is not included in the new code, it should be added here!)

2. witness pass: AA32 = cheby1("LowPass", 4, 1, 32) AND 0.1:0

3. witness adaptation path: AA32 AND AI32 = cheby1("LowPass", 4, 1, 32) AND 0.1:0

4. error path: AA32 AND 0.1:0 AND anti_1Hz. Before I added anti_1Hz filter oaf did nothing. This filter tries to approximate the actuator transfer function. Note, it is not in the witness adaptation path. This is some sort of whitening.

5. correction path: AI32, gain = -1

Convergence time ~ 5 mins. The performance of the filter is far not perfect compared to the offline implementation. But it deals with a stack though.

oaf2.pdf

  6489   Thu Apr 5 07:19:16 2012 kiwamuUpdateLSCDRMI locking

 I tried locking the DRMI to the signal-extraction condition with the new trigger by AS110.

A first thing I tried was : flipping the control sign of the SRCL while keeping the same control setups for the PRCL and MICH.

Occasionally the DRMI was "sort of" locked and hence I believe this setup must be a good starting point.

As a next step I will try some different gains and demodulation phase to make it more lockable.

 


(Time series)

DRMI_2012Apr4_edit.png

 The picture above is time series of some signals when the DRMI was barely locked.
The red arrows indicate the durations when the DRMI was sort of locked.
 (Green curve) REFLDC becoming a high value state, which indicates that the carrier is anti-resonant.
 (Red curve) ASDC becoming dark, which indicates the MICH is in the vicinity of the dark condition.
 (Brown curve) AS110 becoming a high value state, which means the 55 MHz sidebands got amplified by the SRCL.
 (Blue curve) POP22 becoming a high value state, which indicates that the 11 MHz sidebands are resonating in the PRC.
 
According to the measurement of AS110 when PRMI was locked (#6488), the AS110 signal went up to ~ 1 counts or so.
On the other hand when the DRMI was locked the AS110 went to up more than 10 counts as shown in the plot above.
Therefore at least some kind of signal amplification is happening for the 55 MHz sidebands in the SRC.
Looking at the AS CCD, I found that the beam looked like a TEM01 mode (two beam spots at top and bottom) every time when the DRMI was locked.
 
(settings)
  • REFL33I => PRCL  G = -0.2
  • AS55Q => MICH    G = -6
  • AS55I => SRCL     G = 1   (G = -50 for the signal recycling condition)
  • AS55 demod phase = 17 deg
  6488   Thu Apr 5 06:27:51 2012 kiwamuUpdateLSCAS110 sideband monitor installed

[Jenne / Kiwamu]

 We have installed a broad band PD in the AS path in order to monitor the 110 MHz signal associated with the SRC.

The PD is currently connected to the POP110 demodulation board and it seems working fine.

I know this is confusing but right now the signal appears as "POP110" in the LSC front end model.

 


  • Installed a 50% BS at the AS path
    • The AS beam is split to two path - one goes to AS55 and the other goes to the OSA.
    • The new BS is installed on the way of the OSA branch therefore AS55 isn't affected by the new BS.
  • Installed a PDA10A
    • This is a silicon diode with a bandwidth of 150 MHz, and is fast enough to detect the 110 MHz signals.
    • The 110 MHz signal seems going up to approximately -40 dBm according to a coarse measurement with an RF spectrum analyzer.
    • Also a SMA-style high pass filter, HPF-100, was attached to the output to cut off unnecessary sidebands (e.g. 11, 22 MHz and etc.)
  • Put a long BNC cable, which goes from the PD to LSC rack.
    • The end of the cable at the LSC rack was directly connected to the POP110 demod board.
    • The actual POP110 signal path is currently terminated by a 50 Ohm load and therefore this signal  is unavailable.
  • Adjustment of the demodulation phase
    • The demod phase was adjusted to be 7 deg in the EPICS screen. This phase minimize the Q-signal.
    • Locking PRMI with sidebands resonating makes the AS110 signal ~ a few counts and this level is still noticeable.
    • Perhaps we may need to put an RF amplifier to get the signal bigger.
  6487   Thu Apr 5 01:07:08 2012 Mike J.UpdateComputersHysteresis Plots

Here are the hysteresis plots from the most recent model, which uses a modified harmonic oscillator equation y''=-(Frequency)2*y-Hysteresis.  The hysteresis constant seems to change both the amplitude and equilibrium point of the pendulums, which is akin to changing the length of a pendulum without changing the frequency. This does not make sense. Perhaps the hysteresis value should be moved to the "spring" constant for the pendula and not restricted to a position-biasing value.

SHO_hyst_plot.png

  6486   Wed Apr 4 23:57:35 2012 keikoUpdateLSCRAM simulation for Full ifo

 I'm still wondering whether iteration version or simple version is closer approximation to the real situation. Sorry for few explanations here. I will try to present those on Friday.

 

Anyway, here is the results for both:

%*.*.*. Original matrix w/o RAM .*.*.*

REFL f1 : 1.000000        0.000000    -0.000003    -0.000005    0.000007 

  AS f2 : 0.000002        1.000000    0.000009    -0.003522    -0.000002 

 POP f1 : -3954.521443    -0.000965    1.000000    0.019081    -0.000152 

 POP f2 : -32.770726    -0.154433    -0.072594    1.000000    0.024284 

 POP f2 : 922.393978    -0.006608    1.488319    0.042948    1.000000 

 

*** Iteration *** 

%*.*.*. Resulting matrix w/ RAM .*.*.*

REFL f1 : 0.039125    -0.000000    0.003665       0.000005    -0.000007 

  AS f2 : 0.000010    1.000431    0.000009       -0.003500    -0.000002 

 POP f1 : 156.420221    -0.000246    15.586838    0.019406    -0.000154 

 POP f2 : 1.255806    -0.154275    0.047313       1.000008    0.024285 

 POP f2 : -34.814720    -0.006600    -1.884850    0.042950    1.000000 

Offsets converged to:

PRCL =  2.1e-15, MICH = 1.1e-17, SRCL = -3.8e-15, CARM = 2.2e-16, DARM = 0 

(POP CARMs became so much smaller compared with the other matrix below, because the offsets are added al of 5 DoFsl at once here.)

 

*** no iteration, offsets added for each DoF separately ***

REFL f1 : 0.020611        -0.000000    0.003600    0.000005    -0.000007 

AS f2   : 0.000002        1.000000    0.000009    -0.003522    -0.000002 

POP f1  : 1842.776419    -0.000198    21.533358    0.019404    -0.000132 

POP f2  : -32.700639    -0.153095    -0.072481    0.999995    0.024360 

 POP f2 : 922.393862    -0.006435    1.488298    0.042949    0.999982 

Added offsets:

PRCL =  7.5e-15, MICH = 6.25e-16, SRCL = -1.4e-14, CARM = 4.5e-16, DARM = 0

* So far, I used to add all the offsets at once. This time I add CARM and get the CARM row, add PRCL and get the PRCL row... and so on.

Quote:

Koji and Jamie suggested me to include the coupling between DoFs when I calculate the last matrix. So far, I just add all the pos-offsets of 5 DoFs and re-calculate the matrix again. However, once I add one DoF pos-offset, it could already change the LSC matrix therefore different pos-offset to the other four DoF, we must iterate this process until we get the equilibrium pos-offsets for 5 DoFs.

I also noticed an error in the optical configuration file. AM mod levels were smaller than that supposed to be because of the hald power going through the AM-EOMs in the MZI paths. Also I have put PM-Mods in the MZT path which gives the smaller mod indexes. So, smaller mod levels were applied both for PM and AM. As PM-AM ratio is still kept in this, so the matrices were not very wrong, I assume. I'll modify that and post the results again.

 

  6485   Wed Apr 4 21:43:16 2012 Mike J.UpdateComputersBetter Hysteresis Model

A better hysteresis model based on the simple harmonic oscillator equation. Useless variables have been removed and output can now be saved to workspace for plotting. The model is at "/users/mjenson/matlab/SHO_hyst.mdl".

Attachment 1: SHO_hyst.png
SHO_hyst.png
  6484   Wed Apr 4 13:25:29 2012 jamieConfigurationPEMPEM_SLOW (i.e. seismic RMS) channels aquiring

Quote:

I've added the PEM_SLOW.ini file to the fb master file, which should give us the slow seismic RMS channels when the framebuilder is restarted. Example channels:

[C1:PEM-RMS_ACC6_1_3]
[C1:PEM-RMS_GUR2Y_0p3_1]
[C1:PEM-RMS_STS1X_3_10]
etc.

 The framebuilder seems to have been restarted, or restarted on it's own, so these channels are now being acquired.

Below is a minute trend of a smattering of the available RMS channels over the last five days.

2012-04-04-132346_1182x914_scrot.png

  6483   Tue Apr 3 22:50:37 2012 keikoUpdateLSCRAM simulation for Full ifo

Koji and Jamie suggested me to include the coupling between DoFs when I calculate the last matrix. So far, I just add all the pos-offsets of 5 DoFs and re-calculate the matrix again. However, once I add one DoF pos-offset, it could already change the LSC matrix therefore different pos-offset to the other four DoF, we must iterate this process until we get the equilibrium pos-offsets for 5 DoFs.

I also noticed an error in the optical configuration file. AM mod levels were smaller than that supposed to be because of the hald power going through the AM-EOMs in the MZI paths. Also I have put PM-Mods in the MZT path which gives the smaller mod indexes. So, smaller mod levels were applied both for PM and AM. As PM-AM ratio is still kept in this, so the matrices were not very wrong, I assume. I'll modify that and post the results again.

  6482   Tue Apr 3 15:50:58 2012 keikoUpdateLSCRAM simulation for Full ifo

Oops, Yesterday's results for DARM was wrong!

I got more convincing results now. 

 

> (B) Are pos-offsets degrade the CARM and DARM so much (See, the quoted result below), is that true? 

 

Here is the new results. It does change CARM a lot, but not DARM:
 
Matrix1 (normalised so that the diagonals are 1):
REFL f1 : 1.000000    0.000000    0.000008    -0.000005    0.000003 
  AS f2  : 0.000001    1.000000    0.000005    -0.003523    -0.000001 
 POP f1 : -3956.958708    -0.000183    1.000000    0.019064    0.000055 
 POP f2 : -32.766392    -0.154433    -0.072624    1.000000    0.024289 
 POP f2 : 922.415913    -0.006625    1.488912    0.042962    1.000000 
(=Matrix 2)
 
Position offsets:
only CARM, 4.6e-16 (this number changed because I increased the resolution of the calculation)
 
Matrix3 (normalised by matrix 1):
REFL f1 : 0.039780    -0.000000    0.003656    0.000005    -0.000003 
  AS f2  : 0.000008    1.000017    0.000005    -0.003499    -0.000001 
 POP f1 : 159.146819    -0.000138    15.605155    0.019393    0.000055 
 POP f2 : 1.277223    -0.154415    0.047344    1.000008    0.024289 
 POP f2 : -35.422498    -0.006633    -1.886454    0.042963    1.000000 

 

  • CARM got a small position offset which degrades CARM signal 2 orders of mag (still the biggest signal in the sensor, though).
  • DARM was not so bad, and probably the change of the DoF mixture is mostly not changed.
  • Matrices don't change only with 1e-4 RAM. It changes with position offsets.
  • I'll see how the matrix changes without position offsets but only with RAM effects, changing RAM levels.
  • Again, above is C1 configuration, 1e-4 RAM level of PM level.

 

 

Quote:

I add a flow-chart drawing what the scripts do and how the scripts calculate the LSC matrix.

flowchart.png

 

(1) First, you calculate the LSC matrix WITHOUT RAM or anything, just for a reference. This is the first matrix shown in the quoted post.

(2) The script calculates the LSC matrix with RAM. Also, the PDH signals for all 5 DoF are calculated. The PDH signals have offsets due to the RAM effect. The operating position offsets are saved for the next round.

(3) The script calculates the LSC matrix again, with RAM plus the offset of the operation points. The matrix is shown in the last part of the quoted post.

 

Now I am going to check (A) LSC matrices (matrix 2, the second matrix of above chart) with different RAM levels (B) Are pos-offsets degrade the CARM and DARM so much (See, the quated result below), is that true? 

 

  6481   Tue Apr 3 14:17:18 2012 keikoUpdateLSCRAM simulation for Full ifo

I add a flow-chart drawing what the scripts do and how the scripts calculate the LSC matrix.

flowchart.png

 

(1) First, you calculate the LSC matrix WITHOUT RAM or anything, just for a reference. This is the first matrix shown in the quoted post.

(2) The script calculates the LSC matrix with RAM. Also, the heterodyne signals for all 5 DoF are calculated. The signals have offsets due to the RAM effect. The operating position offsets are saved for the next round.

(3) The script calculates the LSC matrix again, with RAM plus the offset of the operation points. The matrix is shown in the last part of the quoted post.

 

Now I am going to check (A) LSC matrices (matrix 2, the second matrix of above chart) with different RAM levels (B) Are pos-offsets degrade the CARM and DARM so much (See, the quated result below), is that true?

Quote:

Original matrix without RAM:

REFL f1 : 1.000000    0.000000    0.000008    -0.000005    0.000003 

  AS f2 : 0.000001    1.000000    0.000005    -0.003523    -0.000001 

 POP f1 : -3956.958708    -0.000183    1.000000    0.019064    0.000055 

 POP f2 : -32.766392    -0.154433    -0.072624    1.000000    0.024289 

 POP f2 : 922.415913    -0.006625    1.488912    0.042962    1.000000 

 

(MICH and SRCL uses the same sensor, with optimised demodulation phase for each DoF.) 

Operation position offsets are:

PRCL   -3.9125e-11 m

SRCL    9.1250e-12 m

CARM  5.0000e-15 m  

and no position offsets for DARM and MICH (because they are differential sensor and not affected by RAM offsets).

 

Resulting matrix with RAM + RAM offsets, normalised by the original matrix:  

REFL f1 : 0.001663    -0.000000    0.003519    0.000005    -0.000003 

  AS f2 : 0.000004    0.514424    0.000004    -0.001676    -0.000001 

 POP f1 : 7.140984    -0.001205    15.051807    0.019254    0.000417 

 POP f2 : 0.029112    -0.319792    0.042583    1.000460    0.024298 

 POP f2 : -0.310318    -0.014385    -1.761519    0.043005    0.999819 

 

 

 

  6480   Tue Apr 3 14:11:33 2012 keikoUpdateLSCRAM simulation for Full ifo

Quote:

Quote:

 I also would like to extend this script to use the DC readout, but don't know how to calculate the postion offset for AS_DC because the error signal is not zero-crossing for AS_DC anymore. Do you have any suggestions for me?

 I don't think I understand the question. AS_DC should not have a zero crossing, correct?

 That's right. I calculate the offset of the operation point (when you have RAM) from the zero-crossing point of the PDH signals. I don't know how to do that for AS_DC, because it doesn't cross zero anymore anytime.

  6479   Tue Apr 3 12:42:19 2012 Mike J.UpdateComputersHysteresis Model

Here's my first hysteresis model in Simulink. It's based on the equation y=Amplitude*sin(frequency*t+phase)+(hysteresis/frequency2) as a solution to y''+frequency2*y+hysteresis=0. All values in the model are variables that should be manipulated through the model workspace or external code.

Attachment 1: hysteresis1.mdl
Model {
  Name			  "hysteresis1"
  Version		  7.6
  MdlSubVersion		  0
  GraphicalInterface {
    NumRootInports	    0
    NumRootOutports	    0
    ParameterArgumentNames  ""
    ComputedModelVersion    "1.9"
    NumModelReferences	    0
... 734 more lines ...
  6478   Tue Apr 3 01:52:15 2012 ZachUpdateLSCRAM simulation for Full ifo

Quote:

 I also would like to extend this script to use the DC readout, but don't know how to calculate the postion offset for AS_DC because the error signal is not zero-crossing for AS_DC anymore. Do you have any suggestions for me?

 I don't think I understand the question. AS_DC should not have a zero crossing, correct?

  6477   Mon Apr 2 23:06:38 2012 SureshUpdateIOOBeam Profile measurement: IPPOS beam: Mystery deepens

Quote:

I fitzed by hand with the numbers for the incident angles on MMT1 and MMT2, and then let the code optimize the position of MMT1 and MMT2.

Here I have set the incident angle for MMT1 = 25deg, and MMT2 = 12deg (should be 3.5deg, 1deg by design).  The length of the telescope doesn't want to change by more than 7mm, but the position of the telescope wants to change by 1.3meters.     Is it possible that the distances on the Monday IPPOS measurements aren't actually correct?

 

I am trying to track down possible errors in our measurements. 

So as a first step I am recomputing the IPPOS waist location with respect to the MC waist, using the same optical layout diagram as the one used by Jenne in her calculations.  Pic of Jenne's lab notebook showing location of optics is attached.  

IPPOS: measurement elog 6459: Vertical Std.Error Horizontal Std.Error
Waist  2.768 mm 5 microns 2.476 mm 10 microns
Waist location from MC waist  12.411 m  17 mm  9.572 m 54 mm

Std Dev of residuals from fit function

  37 microns   54 microns

 Let us compare it with the old measurement of the IPPOS beam from June/18/2010.

IPPOS: measurement June 18th 2010 Vertical Std.Error Horizontal Std.Error
Waist  2. 812mm 8 microns 2.909 mm 20 microns
Waist location from MC waist  9.265 m  224 mm  5.869 m 415 mm

Std Dev of residuals from fit function

  ~ 25 microns   ~25 microns

Note that there is a discrepancy of about 3.2 m in the waist location for the vertical profile and about 3.5 m for the horizontal profile between these two measurements. 

Let us compare these measurements with what is expected from calculations.  Jenne uses the known parameters of MC waist and the locations of the MMT optics to compute the parameters for the IPPOS beam:

IPPOS: Jenne's Calculations elog 6476:

Vertical Std.Error Horizontal Std.Error
Waist  2.844 mm   2.894 mm  
Waist location from MC waist  11.019 m   8.072 m  

 As the 2010 measurements are reported wrt to MMT2 and calculations are wrt MCwaist, I have used the distance between the MCwaist to MMT2 = 3.910 m to shift the reference from MMT2 to MC waist. Refer to the attached diagram from Jenne's notes for this MMT2 <--> MC waist distance.

There is a discrepancy of 1.5 meters between the calculations and recently measured waist location.  The discrepancy with the 18Jun2010 measurement is much larger, about 3 meters in both v and h.

Are such variations to be expected between two successive measurements?  I looked at another case where we have two measurements of a beam to see what to expect.

I looked at the REFL (Reflection from PRM) case, where we repeated a measurement, to see how much variation could happen in w0 and zr, between repeated measurements. This was a particularly bad case as our first attempt had problems due to OL servo loop oscillations in the PRM suspension damping.  We fixed that later and measurement 2 has smaller residuals. And I think we are doing okay in IPPOS case as seen by the reduced scatter of the residuals.

 These are the fits from the REFL beam measurement 1

REFL: Reflection from PRM: measurement 1 Vertical Error Horizontal Error
Waist 1.662 mm 4 microns 2.185 mm 4 microns
Waist location from MMT2 after reflection at PRM 1.781 m 17 mm 4.620 m 53 mm
Std.Dev. of residuals from fit function   61 microns   98 microns

 I have also recomputed the fits to the data from REFL beam measurement 2.  They match the earlier fits reported by kiwamu in his elog 6446

REFL: Reflection from PRM: measurement 2 Vertical Error Horizontal Error
Waist 1.511 mm 3 microns 2.128 mm 3 microns
Waist location from MMT2 after reflection at PRM 1.281 m 9 mm 3.211 m 37 mm
Std. Dev of residuals from fit function   58 microns   61 microns

 

Note that between these two measurements the beam waist location has shifted by 0.5 m for the vertical and about 1.3 m for the horizontal cases.  So variations of 1.5 m in the waist locations are possible if we are not careful.  But this is a particularly extreme example, I think we are doing better now and the measurement is unlikely to change significantly if we repeat it.

Some notes:

Fits for IPPOS and both REFL measurements 1 and 2 are attached. 

The zero reference for the z axis of the IPPOS beam plot is at a distance of 6.719 m from MC waist for a beam propagating towards the IPPOS QPD.

The zero reference for the z axis of the REFL beam plots is at a distance of 5.741 m from the MMT2 in the direction of a beam reflected by PRM and propagating towards the REFL port.

 

Attachment 1: 40mOpticsLocations.pdf
40mOpticsLocations.pdf
Attachment 2: Beam-Profile_IPPOS_wError.pdf
Beam-Profile_IPPOS_wError.pdf
Attachment 3: Beam-Profile_PRM_1_wError.pdf
Beam-Profile_PRM_1_wError.pdf
  6476   Mon Apr 2 18:58:32 2012 JenneUpdateIOOBeam Profile measurement: IPPOS beam

Suresh noted that I never wrote down the waist positions of the beam propagated through the MMT (based on where we think it is from ruler-based measurements).  This uses the MC waist measured beam as the starting point, and just propagates through the MMT, out to the IPPOS port.

Yq:

  Properties:
                    q: 2.2488 +23.8777i
               lambda: 1.0640e-06
            waistSize: 0.0028  (at z = 13.2676 meters)
               waistZ: 2.2488    (relative to z = 13.2676 meters)
      divergenceAngle: 1.1910e-04
    radiusOfCurvature: 255.7849
            beamWidth: 0.0029
        rayleighRange: 23.8777

So, to sum up, the vertical beam waist is 2.8438 mm at 11.0188 meters from the MC waist.

 

Xq:

  Properties:
                    q: 5.1953 +24.7342i
               lambda: 1.0640e-06
            waistSize: 0.0029   (at z = 13.2676 meters)
               waistZ: 5.1953    (relative to z = 13.2676 meters)
      divergenceAngle: 1.1702e-04
    radiusOfCurvature: 122.9525
            beamWidth: 0.0030
        rayleighRange: 24.7342

So, to sum up, the horizontal beam waist is 2.8943 mm at 8.0723 meters from the MC waist.

In pictorial form,

NonOptimized_propagatedWaistPlotted_LowRes.png

  6475   Mon Apr 2 18:24:34 2012 keikoUpdateLSCRAM simulation for Full ifo

 I extended my RAM script from DRMI (3DoF) to the full IFO (5DoF).

Again, it calculates the operation point offsets for each DoF from the opt model with RAM. Then the position offsets are added to the model, and calculates the LSC matrix. RAM level is assumed as 0.1% of the PM modulation level, as usual, and lossless for a simple model.

 

 

Original matrix without RAM:

REFL f1 : 1.000000    0.000000    0.000008    -0.000005    0.000003 

  AS f2 : 0.000001    1.000000    0.000005    -0.003523    -0.000001 

 POP f1 : -3956.958708    -0.000183    1.000000    0.019064    0.000055 

 POP f2 : -32.766392    -0.154433    -0.072624    1.000000    0.024289 

 POP f2 : 922.415913    -0.006625    1.488912    0.042962    1.000000 

 

(MICH and SRCL uses the same sensor, with optimised demodulation phase for each DoF.) 

Operation position offsets are:

PRCL   -3.9125e-11 m

SRCL    9.1250e-12 m

CARM  5.0000e-15 m  

and no position offsets for DARM and MICH (because they are differential sensor and not affected by RAM offsets).

 

Resulting matrix with RAM + RAM offsets, normalised by the original matrix:  

REFL f1 : 0.001663    -0.000000    0.003519    0.000005    -0.000003 

  AS f2 : 0.000004    0.514424    0.000004    -0.001676    -0.000001 

 POP f1 : 7.140984    -0.001205    15.051807    0.019254    0.000417 

 POP f2 : 0.029112    -0.319792    0.042583    1.000460    0.024298 

 POP f2 : -0.310318    -0.014385    -1.761519    0.043005    0.999819 

 

As you can see in the second matrix, the CARM and DARM rows are completely destroyed by the RAM offsets! The signals are half reduced in the DARM case, so the mixture between DARM and MICH are about 50% degraded.

 I also would like to extend this script to use the DC readout, but don't know how to calculate the postion offset for AS_DC because the error signal is not zero-crossing for AS_DC anymore. Do you have any suggestions for me?

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

  6474   Sat Mar 31 08:01:07 2012 kiwamuUpdateLSCDRMI measurement

I have measured the sensing matrix of the DRMI although the lock still doesn't stay for a long time.

As for the noise budget, it looks very tough as there are more glitches than that in the PRMI.

In this weekend I will take some more trials in the DRMI lock until I am satisfied.

  6473   Fri Mar 30 17:37:09 2012 steveUpdateGeneralcutting green welding glass for beam dumps

Quote:

Schott, green welding glass, shade 14, 3 mm thick  was measured in the beam path of 1.2W, S polarization of 1064nm at ~1 mm diameter size as MC reflected path.

Absorption 95%, R 5% at incident angle 25-50 degrees. It looks like the perfect material for beam trap.

 

 The CIT Chemistry Glass Shop cuts turned out to be sloppy using diamond disc blade  cutter.

East coast Precision Glass & Optics  offered scribed cut and polished side. Their quote price was high and time consuming.

The GLASS HOUSE  shop in Pasadena 626 / 796-9151 on Walnut did a good and cheap job. Oscar the cutter will finish the rest of the cutting by next Friday, April 6

He used scribed cutting technic and his 1" x  0.5"  pieces are good. Bob will pick up them up.

Attachment 1: IMG_0096.JPG
IMG_0096.JPG
ELOG V3.1.3-